OpenStudy (anonymous):

(3x)^2+wz-6x= Evaluate. w = 3, x = 6, y = 5, and z = 9

OpenStudy (anonymous):

(3*6)^2+3*9-6*6=(18)^2+27-36=324+27-36=351-36=315 value of y is given but have no use here...

OpenStudy (anonymous):

so its 315

OpenStudy (anonymous):

yes...